OpenStudy (anonymous):

y'' - 4y' + 4y = 3 δ ( t - 1) + δ ( t-2) by solving Laplace. help?

OpenStudy (anonymous):

This looks like Satan's work.

OpenStudy (usukidoll):

lol

OpenStudy (usukidoll):

I'm not there yet but I heard it is nastayt

OpenStudy (anonymous):

i understand the first part but the last one is kinda weird. What section is this on Laplace?

OpenStudy (anonymous):

the initial values it is y(0) = 1 and y'(0) = 1

OpenStudy (usukidoll):

is it one of those subsitutue forms?

OpenStudy (anonymous):

it is function delta for Dirac

OpenStudy (loser66):

@ybarrap

OpenStudy (ybarrap):

The Laplace of the delta functions on the right side of the equal sign is: $$ \Large{ e^{-2 s}+3 e^{-s} } $$ Combining with Laplace on the left: $$ \Large{ s^2 L_t[y(t)](s))+4 L_t[y(t)](s)\\ \qquad -4 (s (L_t[y(t)](s))-y(0))-s y(0)-y'(0)\\ =s^2Y(s)+4Y(s)-4sY(s)-y(0)-sy(0)-y'(0)\\ =Y(s)(s^2-4s+4)-y(0)(1-s)-y'(0)\\ =Y(s)(s^2-4s+4)-1(1-s)-1\\ =Y(s)(s^2-4s+4)-2+s=e^{-2 s}+3 e^{-s}\\ Y(s)=\cfrac{e^{-2 s}+3 e^{-s}+2-s}{s^2-4s+4} } $$ Find the inverse of this to get y(t).

OpenStudy (loser66):

@ybarrap you made mistake at 4L'(y) it is 4SL(y) - 4 , not just -1

OpenStudy (anonymous):

exact

OpenStudy (loser66):

the numerator is s -3 ( the last term) not -s +2

OpenStudy (loser66):

however, it's not big deal, next step is most important

OpenStudy (ybarrap):

you're right, but this is the basic idea. Hopefully it helps :)

OpenStudy (usukidoll):

@ybarrap can you help me with my question after this?

OpenStudy (ybarrap):

sure

OpenStudy (usukidoll):

THANK GAWD!

OpenStudy (loser66):

@ybarrap not finish yet, Socrates' friend.

OpenStudy (usukidoll):

D:

OpenStudy (usukidoll):

nooo!!!!! ((cries**

OpenStudy (ybarrap):

Just saw your work there, quite involved @UsukDoll

OpenStudy (usukidoll):

but but...I just ughhhhhh I know I have to work backwards but man I'm stuck as the hexagon hills of the eternal realm darkness stuff

OpenStudy (ybarrap):

The inverse here will also take some work - @Loser66 is an expert at Laplace inverses

OpenStudy (ybarrap):

She can get the job done quicker than everyone on this site

OpenStudy (usukidoll):

please save my differential equations soul. x(

OpenStudy (anonymous):

please. stop

OpenStudy (ybarrap):

@marielen123 - Does the work here make sense so far?

OpenStudy (anonymous):

No

OpenStudy (ybarrap):

The Laplace transform part or reducing it after the transform?

OpenStudy (anonymous):

reducing it after the transform,

OpenStudy (anonymous):

y(s) = [3e^-s + e ^-2s + s - 3] / s^2 -4s + 4

OpenStudy (anonymous):

do not think the answer table

OpenStudy (loser66):

how? without looking at the table, how can you find out the inverse of them?

OpenStudy (usukidoll):

you need the tables...

OpenStudy (anonymous):

I could settle for partial fraction or looking table

OpenStudy (anonymous):

but I do not see

OpenStudy (loser66):

sadly, I have to say: no way. even you can partial fraction the denominator, you don't have formula for the numerator one. I mean for \(\dfrac{3e^{-3s}}{s^2-4s+4}\) after doing partial, what do you have? you cannot because they don't have the same variables or coefficient

OpenStudy (ybarrap):

Take the 1st part, the inverse of y(s) = [3e^-s ] / s^2 -4s + 4 is $$ \Large{ 3 e^{2 (t-1)} (t-1) u(t-1) } $$ u(t) is the unit step function. Note that s^2 - 4s + 4 factors to (s-2)^2 in the denominator. Then use rules for time shifting. You can use tables or wolfram: http://www.wolframalpha.com/input/?i=inverse+laplace+transform+of++ [3e^-s+]+%2F+%28s^2+-4s+%2B+4+%29

OpenStudy (ybarrap):

When you partial expand this1st term you get: $$ \Large {\\ (3 e^{-s} s)/(s-2)^2-(3 e^{-s})/(s-2)-(3 e^{-s})/(s-2)^2 } $$

OpenStudy (anonymous):

I did no understand

OpenStudy (ybarrap):

Do you know who to use partial fraction expansion? This is critical.

OpenStudy (usukidoll):

oh I remember that calculus ii stuff

OpenStudy (usukidoll):

when the bottom is factorable you have something like A B C on the numerator but if the equation on the denominator can't be factorable, you have something like Ax+B

OpenStudy (ybarrap):

Here is a quick refresher: http://en.wikipedia.org/wiki/Partial_fraction_decomposition

OpenStudy (anonymous):

I know A/ (s-2)^2 + B / (s-2) ^2

OpenStudy (usukidoll):

nooooooooooooooooooooooo you need to split that up

OpenStudy (usukidoll):